A variable, X, is uniformly distributed between 10 and 24.

(a) What is Pr(16<X<20)?

(b) What is the mean and variance of X?

(c) What is the 75th percentile for this variable?
Answer #1

Solution :

Given that,

a = 10

b = 24

P(c < x < d) = (d - c) / (b - a)

P(16< x < 20) = (20- 16) / (24 - 10) = 0.2857

Probability = 0.2857

b.

variance =  (b - a)2 / 12

= (24 - 10)2 / 12

=16.33

mean = (a + b) / 2=24+10/2=17

c.

PERCENTILE

= a + (b - a)p

=24+ (24 - 10)0.75

=34.5